# Quasinormal modes of hairy black holes in shift-symmetric theories

### v1

@@ Abstract
Alternative theories of gravity may be put to the test by making use of gravitational wave observations. Scalar-tensor theories provide a relatively simple framework that allows for compact object solutions with characteristics different from those of their GR counterparts. In shift-symmetric theories, black hole hair may arise when a linear coupling with the Gauss-Bonnet (GB) invariant is introduced. The effect of this coupling on the black hole quasinormal modes (QNMs) has been shown to be small, therefore limiting the observational interest in the theory. In general, however, one expects that additional shift-symmetric terms in the Lagrangian may be relevant, and indeed, it has been shown that they can have a nontrivial impact on the mass and scalar charge of the stationary solutions. In this work we explore the effect that various such terms have on the axial and polar QNMs of the hairy black hole solutions.
@@ II.1 Shift-symmetric Horndeski gravity
$\mathcal{L}_{5}=$ $\,G_{5}(X)G_{\mu\nu}\nabla^{\mu}\nabla^{\nu}\phi-\frac{G_{5X}}{6}\big{[}\left(\Box\phi\right)^{3}-3\Box\phi(\nabla_{\mu}\nabla_{\nu}\phi)^{2}+2(\nabla_{\mu}\nabla_{\nu}\phi)^{3}\big{]}\,,$ (5)
@@
$S=\frac{1}{2k}\int\mathrm{d}^{4}x\,\sqrt{-g}\bigg{[}\frac{R}{2}+X+\alpha\,\phi\,\mathcal{G}+\gamma\,G_{\mu\nu}\nabla^{\mu}\phi\nabla^{\nu}\phi+\sigma X\Box\phi+\kappa\,X^{2}\bigg{]}.$ (9)
@@ II.2 Background solutions
II: $\sqrt{r_{h}^{6}-576\alpha^{2}\gamma-24\alpha(8\alpha+\sigma)r_{h}^{2}}\times\big{[}24\alpha\gamma r_{h}+(4\alpha+\sigma)r_{h}^{3}\big{]}+4\alpha\big{[}r_{h}^{6}-576\alpha^{2}\gamma-24\alpha(8\alpha+\sigma)r_{h}^{2}\big{]}\neq 0\,.$ (12)
@@ III.1 Axial modes
$g(r)=\frac{dr_{*}}{dr}\quad\text{with}\quad g^{2}(r)=\frac{-4\alpha B^{\prime}\phi^{\prime}-B(8\alpha\phi^{\prime\prime}+\gamma\phi^{\prime 2})+1}{B(-4\alpha BA^{\prime}\phi^{\prime}+AB\gamma\phi^{\prime 2}+A)}\,,$ (25)
@@ III.2 Polar modes
${\textbf{X}}_{p}=\big{(}\boldsymbol{\Psi}^{p}_{(h,1)},\boldsymbol{\Psi}^{p}_{(h,2)},\boldsymbol{\Psi}^{p}_{(\infty,1)},\boldsymbol{\Psi}^{p}_{(\infty,2)}\big{)}=\begin{bmatrix}K^{(h,1)}&K^{(h,2)}&K^{(\infty,1)}&K^{(\infty,2)}\\ H_{1}^{(h,1)}&H_{1}^{(h,1)}&H_{1}^{(\infty,1)}&H_{1}^{(\infty,2)}\\ \phi_{1}^{(h,1)}&\phi_{1}^{(h,1)}&\phi_{1}^{(\infty,1)}&\phi_{1}^{(\infty,2)}\\ \phi_{1}^{\prime(h,1)}&\phi_{1}^{\prime(h,1)}&\phi_{1}^{\prime(\infty,1)}&\phi_{1}^{\prime(\infty,2)}\\ \end{bmatrix}\,,$ (40)
@@
Polar gravitational-led modes for $\boldsymbol{\ell=2}$
Polar scalar-led modes for $\boldsymbol{\ell=2}$
@@
Before closing this section, let us comment on differences on the QNMs arising in the various scenarios for same-mass black holes in different scenarios. A reasonable question to ask is the following: among the three additional shift-symmetric terms we considered in this work, is there any one in particular that is expected to yield more important contributions than the rest? As it became clear from the analysis of both the axial and polar spectrum, the $\kappa$-term has very limited effects on the QNMs in comparison with $\gamma$ and $\sigma$. This result was expected based on the effect that same term has on the scalar charge of the stationary solutions. Between $\gamma$ and $\sigma$ the effects were comparable towards the edge of the existence lines, but $\gamma$ seemed to exert a stronger effect on the QNMs for heavier black holes (i.e. for smaller $\alpha/M^{2}$). In Tab. 1 and 2 we present the results for the QNMs in four different models characterized by the parameters $(\tilde{\gamma},\tilde{\sigma},\tilde{\kappa})$, with $\ell=2$ and $\alpha/M^{2}=0.2$. The four different cases are supplemented with the simplest case where all the additional shift-symmetric couplings are ignored, which is shown in the first row of the tables. The fourth case we considered corresponds to an $\mathcal{O}(1)$ combination of all three terms. Notice that in order to be able to compare same-mass black holes between the different models we have to resort to a value for $\alpha/M^{2}$ which yields hairy solutions for all models simultaneously. We have fixed that value to be $0.2$. From the table we see that $\gamma<0$ appears to be the most influential among the four scenarios, even considering the mixed one with $(\tilde{\gamma},\tilde{\sigma},\tilde{\kappa})=(-1,-1,-1)$. This further strengthens the conclusion that $\gamma$ is the most dominant of the terms we considered.
@@ IV Astrophysical implications
$S=\int d^{4}x\sqrt{-g}\left[\frac{R}{2k}+\bar{X}+\bar{\alpha}\varphi\,\mathscr{G}\right]=\frac{1}{k}\int d^{4}x\sqrt{-g}\left[\frac{R}{2}+k\bar{X}+(\sqrt{k}\bar{\alpha})(\sqrt{k}\varphi)\,\mathscr{G}\right]\,,$ (41)
@@ V Conclusions
For the polar modes the analysis presents extra complexity since now two types of modes can be found, namely gravitational-led and scalar-led ones. As expected, the isospectrality between axial and polar modes was broken. Similarly to the axial case we found the $\gamma$ and $\sigma$ terms to be the ones that predominantly affect the modes. Especially in the scalar-led modes the effects were very important both in the real and imaginary part of the frequency. One interesting characteristic that we observed relates to the fact that by including the $\sigma$ and $\gamma$ contributions, the imaginary part of the frequency was in general larger in the polar modes and not in the axial ones, which is in contrast to the case when we only consider the $\phi\mathscr{G}$ interaction.
@@
Our results showed that the frequencies and decay rates of the modes may vary considerably depending on the exact model. It would, therefore, be important to examine in detail the ability of future interferometers to access these effects. At the same time, a thorough analysis of potential new constraints on the parameters of the theory should be performed. Finally, there is a need for an extensive analysis of the exact region of the parameter space where a perturbative approach in accordance with D’Addario et al. (2023) is possible. This will allow us to conclusively determine whether or not results such as the ones presented in the current work will be relevant for LISA. These tasks go beyond the purpose of this paper and are left for future work.
@@ Acknowledgements.
Appendix A Background and perturbation equations
@@
$G_{\mu\nu}-\nabla_{\mu}\phi\nabla_{\nu}\phi-\frac{1}{2}g_{\mu\nu}(\nabla\phi)^{2}+\frac{\alpha}{g}g_{\mu(\rho}g_{\sigma)\nu}\epsilon^{\kappa\rho\alpha\beta}\epsilon^{\sigma\gamma\lambda\tau}{R}_{\lambda\tau\alpha\beta}\nabla_{\gamma}\nabla_{\kappa}\phi$
$-\gamma\big{[}2\nabla_{\nu}\nabla_{\mu}\phi\,\Box\phi-{R}\nabla_{\mu}\phi\nabla_{\nu}\phi-2G_{\nu\rho}\nabla_{\mu}\phi\nabla^{\rho}\phi-2G_{\mu\rho}\nabla_{\nu}\phi\nabla^{\rho}\phi+{R}_{\mu\nu}(\nabla_{\mu}\phi)^{2}-2\nabla_{\rho}\nabla_{\nu}\phi\nabla^{\rho}\nabla_{\mu}\phi-g_{\mu\nu}(\Box\phi)^{2}$ (42)
$+G_{\rho\sigma}g_{\mu\nu}\nabla^{\rho}\phi\nabla^{\sigma}\phi+g_{\mu\nu}{R}_{\rho\sigma}\nabla^{\rho}\phi\nabla^{\sigma}\phi-2{R}_{\mu\rho\nu\sigma}\nabla^{\rho}\phi\nabla^{\sigma}\phi+g_{\mu\nu}(\nabla_{\sigma}\nabla_{\rho}\phi)^{2}\big{]}$
$-\sigma\big{[}\nabla_{\mu}\phi\nabla_{\nu}\phi\;\Box\phi-2\nabla_{\rho}\nabla_{(\mu}\phi\nabla_{\nu)}\phi\nabla^{\rho}\phi+g_{\mu\nu}\nabla^{\rho}\phi\nabla_{\sigma}\nabla_{\rho}\phi\nabla^{\sigma}\phi\big{]}+\kappa\big{[}g_{\mu\nu}(\nabla\phi)^{2}(\nabla\phi)^{2}/4-(\nabla\phi)^{2}\nabla_{\mu}\phi\nabla_{\nu}\phi\big{]}=0\,,$
while the scalar one by:
$\Box\phi+\alpha\mathscr{G}+2\gamma G^{\mu\nu}\nabla_{\nu}\nabla_{\mu}\phi-\kappa\big{[}(\nabla\phi)^{2}\Box\phi+2\nabla^{\mu}\phi\nabla^{\nu}\phi\nabla_{\mu}\nabla_{\nu}\phi\big{]}-\sigma\big{[}\nabla^{\mu}\phi\,\Box\,\nabla_{\mu}\phi+(\nabla_{\mu}\nabla_{\nu}\phi)^{2}-$ (43)
$\nabla_{\mu}\,\Box\phi\nabla^{\mu}\phi-(\Box\phi)^{2}\big{]}=0\,.$
@@
$(tt):\quad$ $2B(-2(8\alpha\phi^{\prime\prime}+1)+B^{\prime}\phi^{\prime}(24\alpha+r^{2}\sigma\phi^{\prime 2}-6\gamma r\phi^{\prime})-((2\gamma+r^{2})(\phi^{\prime})^{2}))-4B^{\prime}(4\alpha\phi^{\prime}+r)$ (44)
$+B^{2}(32\alpha\phi^{\prime\prime}-4\phi^{\prime 2}(\gamma-r^{2}\sigma\phi^{\prime\prime})+\kappa r^{2}\phi^{\prime 4}-16\gamma r\phi^{\prime}\phi^{\prime\prime})+4=0\,,$
$(rr):\quad$ $2A^{\prime}B(r^{2}\sigma B\phi^{\prime 3}+8\alpha(3B-1)\phi^{\prime}-6\gamma rB\phi^{\prime 2}-2r)+A(B^{2}\phi^{\prime 2}(-12\gamma-3\kappa r^{2}\phi^{\prime 2}+8r\sigma\phi^{\prime})$ (45)
$+2B((2\gamma+r^{2})\phi^{\prime 2}-2)+4)=0\,,$
(ϕ)::italic-$(\phi):\quad$ $2A(\phi^{\prime}(4Br\left(-B\gamma A^{\prime\prime}+2AB\sigma\phi^{\prime\prime}+A\right)+AB^{\prime}(-6B\gamma+2\gamma+r^{2}))+B\phi^{\prime 2}(B(r^{2}\sigma A^{\prime\prime}-6A\kappa r^{2}\phi^{\prime\prime}$
$+4A\sigma)+6Ar\sigma B^{\prime})+2B(4\alpha(B-1)A^{\prime\prime}+A\phi^{\prime\prime}(-2B\gamma+2\gamma+r^{2}))-AB\kappa r\phi^{\prime 3}(3rB^{\prime}+4B))$ (46)
$+AA^{\prime}(B^{\prime}(8\alpha(3B-1)+3Br^{2}\sigma\phi^{\prime 2}-12B\gamma r\phi^{\prime})+2B(\phi^{\prime}(-6B\gamma+2Br^{2}\sigma\phi^{\prime\prime}+2\gamma+r^{2})$
$-B\kappa r^{2}\phi^{\prime 3}-4B\gamma r\phi^{\prime\prime}+4Br\sigma\phi^{\prime 2}))+BA^{\prime 2}(-8\alpha(B-1)-Br^{2}\sigma\phi^{\prime 2}+4B\gamma r\phi^{\prime})=0\,.$
@@
Next, we present the form of the perturbation equations in the frequency domain. These are the equations that we integrate in our search for the QNMs. In the following equations the functions $\mathcal{F}_{i}$ depend on the background solutions and on the frequency $\omega$. They are rather complicated so we avoid providing their exact form here, but we denote the parameters of the theory upon which they depend.

### v2

@@ Abstract
Alternative theories of gravity may be put to the test by making use of gravitational wave observations. Scalar-tensor theories provide a relatively simple framework that allows for compact object solutions with characteristics different from those of their general relativity counterparts. In shift-symmetric theories, black hole hair may arise when a linear coupling with the Gauss-Bonnet (GB) invariant is introduced. The effect of this coupling on the black hole quasinormal modes (QNMs) has been shown to be small, therefore limiting the observational interest in the theory. In general, however, one expects that additional shift-symmetric terms in the Lagrangian may be relevant, and indeed, it has been shown that they can have a nontrivial impact on the mass and scalar charge of the stationary solutions. In this work we explore the effect that various such terms have on the axial and polar QNMs of the hairy black hole solutions.
@@ II.1 Shift-symmetric Horndeski gravity
$\mathcal{L}_{5}=$ $\,G_{5}(X)G_{\mu\nu}\nabla^{\mu}\nabla^{\nu}\phi-\frac{G_{5X}}{6}\big{[}\left(\Box\phi\right)^{3}$ (5)
$-3\Box\phi(\nabla_{\mu}\nabla_{\nu}\phi)^{2}+2(\nabla_{\mu}\nabla_{\nu}\phi)^{3}\big{]}\,,$ (6)
@@
$\begin{split}S=\frac{1}{2k}\int\mathrm{d}^{4}&x\,\sqrt{-g}\bigg{[}\frac{R}{2}+X+\alpha\,\phi\,\mathcal{G}\\ &+\gamma\,G_{\mu\nu}\nabla^{\mu}\phi\nabla^{\nu}\phi+\sigma X\Box\phi+\kappa\,X^{2}\bigg{]}.\end{split}$ (10)
@@ II.2 Background solutions
$\begin{split}\textbf{II:}\quad&\sqrt{r_{h}^{6}-576\alpha^{2}\gamma-24\alpha(8\alpha+\sigma)r_{h}^{2}}\\ &\times\big{[}24\alpha\gamma r_{h}+(4\alpha+\sigma)r_{h}^{3}\big{]}+4\alpha\big{[}r_{h}^{6}-576\alpha^{2}\gamma\\ &\qquad-24\alpha(8\alpha+\sigma)r_{h}^{2}\big{]}\neq 0\,.\end{split}$ (13)
@@ III.1 Axial modes
$\frac{dr_{*}}{dr}\equiv g(r)=\sqrt{\frac{1-4\alpha B^{\prime}\phi^{\prime}-B(8\alpha\phi^{\prime\prime}+\gamma\phi^{\prime 2})}{B(-4\alpha BA^{\prime}\phi^{\prime}+AB\gamma\phi^{\prime 2}+A)}}\,.$ (26)
@@ III.2 Polar modes
Polar gravitational-led modes for $\boldsymbol{\ell=2}$
Polar scalar-led modes for $\boldsymbol{\ell=2}$
@@
$\begin{split}{\textbf{X}}_{p}=&\;\big{(}\boldsymbol{\Psi}^{p}_{(h,1)},\boldsymbol{\Psi}^{p}_{(h,2)},\boldsymbol{\Psi}^{p}_{(\infty,1)},\boldsymbol{\Psi}^{p}_{(\infty,2)}\big{)}\\[5.69054pt] =&\begin{bmatrix}K^{(h,1)}&K^{(h,2)}&K^{(\infty,1)}&K^{(\infty,2)}\\ H_{1}^{(h,1)}&H_{1}^{(h,1)}&H_{1}^{(\infty,1)}&H_{1}^{(\infty,2)}\\ \phi_{1}^{(h,1)}&\phi_{1}^{(h,1)}&\phi_{1}^{(\infty,1)}&\phi_{1}^{(\infty,2)}\\ \phi_{1}^{\prime(h,1)}&\phi_{1}^{\prime(h,1)}&\phi_{1}^{\prime(\infty,1)}&\phi_{1}^{\prime(\infty,2)}\\ \end{bmatrix}\,,\end{split}$ (41)
@@
Before closing this section, let us comment on differences on the QNMs arising in the various scenarios for same-mass black holes in different scenarios. A reasonable question to ask is the following: among the three additional shift-symmetric terms we considered in this work, is there any one in particular that is expected to yield more important contributions than the rest? As it became clear from the analysis of both the axial and polar spectrum, the $\kappa$-term has very limited effects on the QNMs in comparison with $\gamma$ and $\sigma$. This result was expected based on the effect that same term has on the scalar charge of the stationary solutions. Between $\gamma$ and $\sigma$ the effects were comparable towards the edge of the existence lines, but $\gamma$ seemed to exert a stronger effect on the QNMs for heavier black holes (i.e. for smaller $\alpha/M^{2}$). Since larger mass black holes are particularly interesting from an observation point of view (see Sec. IV), the contribution of the $\gamma$ term is expected to be the most interesting in comparison with $\sigma$ and $\kappa$. In Tab. 1 and 2 we present the results for the QNMs in four different models characterized by the parameters $(\tilde{\gamma},\tilde{\sigma},\tilde{\kappa})$, with $\ell=2$ and $\alpha/M^{2}=0.2$. The four different cases are supplemented with the simplest case where all the additional shift-symmetric couplings are ignored, which is shown in the first row of the tables. The fourth case we considered corresponds to an $\mathcal{O}(1)$ combination of all three terms. Notice that in order to be able to compare same-mass black holes between the different models we have to resort to a value for $\alpha/M^{2}$ which yields hairy solutions for all models simultaneously. We have fixed that value to be $0.2$. From the table we see that $\gamma<0$ appears to be the most influential among the four scenarios, even considering the mixed one with $(\tilde{\gamma},\tilde{\sigma},\tilde{\kappa})=(-1,-1,-1)$. This further strengthens the conclusion that $\gamma$ is the most dominant of the terms we considered.
@@ IV Astrophysical implications
$\begin{split}S=&\int d^{4}x\sqrt{-g}\left[\frac{R}{2k}+\bar{X}+\bar{\alpha}\varphi\,\mathscr{G}\right]=\\ &\frac{1}{k}\int d^{4}x\sqrt{-g}\left[\frac{R}{2}+k\bar{X}+(\sqrt{k}\bar{\alpha})(\sqrt{k}\varphi)\,\mathscr{G}\right]\,,\end{split}$ (42)
@@ V Conclusions
For the polar modes the analysis presents extra complexity since now two types of modes can be found, namely gravitational-led and scalar-led ones. The isospectrality between axial and polar modes was broken. Similarly to the axial case we found the $\gamma$ and $\sigma$ terms to be the ones that predominantly affect the modes. Especially in the scalar-led modes the effects were very important both in the real and imaginary part of the frequency. One interesting characteristic that we observed relates to the fact that by including the $\sigma$ and $\gamma$ contributions, the imaginary part of the frequency was in general larger in the polar modes and not in the axial ones, which is in contrast to the case when we only consider the $\phi\mathscr{G}$ interaction.
@@
Our results showed that the frequencies and decay rates of the modes may vary considerably depending on the exact model. It would, therefore, be important to examine in detail the ability of future interferometers to access these effects. At the same time, a thorough analysis of potential new constraints on the parameters of the theory should be performed. Finally, there is a need for an extensive analysis of the exact region of the parameter space where a perturbative approach in accordance with D’Addario et al. (2023) is possible. This will allow us to conclusively determine whether or not results such as the ones presented in the current work will be relevant for LISA. These tasks go beyond the purpose of this paper and are left for future work. Finally, a potential continuation of the current work could examine the grey-body factors in generic frameworks such as the shift-symmetric one studied here.
@@ Acknowledgements.
Appendix A Background equations
@@
$G_{\mu\nu}-\nabla_{\mu}\phi\nabla_{\nu}\phi-\frac{1}{2}g_{\mu\nu}(\nabla\phi)^{2}$
$+\frac{\alpha}{g}g_{\mu(\rho}g_{\sigma)\nu}\epsilon^{\kappa\rho\alpha\beta}\epsilon^{\sigma\gamma\lambda\tau}{R}_{\lambda\tau\alpha\beta}\nabla_{\gamma}\nabla_{\kappa}\phi$
$-\gamma\big{[}2\nabla_{\nu}\nabla_{\mu}\phi\,\Box\phi-{R}\nabla_{\mu}\phi\nabla_{\nu}\phi-2G_{\nu\rho}\nabla_{\mu}\phi\nabla^{\rho}\phi$
$-2G_{\mu\rho}\nabla_{\nu}\phi\nabla^{\rho}\phi+{R}_{\mu\nu}(\nabla_{\mu}\phi)^{2}$ (43)
$-2\nabla_{\rho}\nabla_{\nu}\phi\nabla^{\rho}\nabla_{\mu}\phi-g_{\mu\nu}(\Box\phi)^{2}+G_{\rho\sigma}g_{\mu\nu}\nabla^{\rho}\phi\nabla^{\sigma}\phi$
$+g_{\mu\nu}{R}_{\rho\sigma}\nabla^{\rho}\phi\nabla^{\sigma}\phi-2{R}_{\mu\rho\nu\sigma}\nabla^{\rho}\phi\nabla^{\sigma}\phi+g_{\mu\nu}(\nabla_{\sigma}\nabla_{\rho}\phi)^{2}\big{]}$
$-\sigma\big{[}\nabla_{\mu}\phi\nabla_{\nu}\phi\;\Box\phi-2\nabla_{\rho}\nabla_{(\mu}\phi\nabla_{\nu)}\phi\nabla^{\rho}\phi$
$+g_{\mu\nu}\nabla^{\rho}\phi\nabla_{\sigma}\nabla_{\rho}\phi\nabla^{\sigma}\phi\big{]}+\kappa\big{[}g_{\mu\nu}(\nabla\phi)^{2}(\nabla\phi)^{2}/4$
$-(\nabla\phi)^{2}\nabla_{\mu}\phi\nabla_{\nu}\phi\big{]}=0\,,$
while the scalar one is given by:
$\Box\phi+\alpha\mathscr{G}+2\gamma G^{\mu\nu}\nabla_{\nu}\nabla_{\mu}\phi-\kappa\big{[}(\nabla\phi)^{2}\Box\phi$
$+2\nabla^{\mu}\phi\nabla^{\nu}\phi\nabla_{\mu}\nabla_{\nu}\phi\big{]}-\sigma\big{[}\nabla^{\mu}\phi\,\Box\,\nabla_{\mu}\phi$ (44)
$+(\nabla_{\mu}\nabla_{\nu}\phi)^{2}-\nabla_{\mu}\,\Box\phi\nabla^{\mu}\phi-(\Box\phi)^{2}\big{]}=0\,.$
@@
$(tt):\quad$ $2B(-2(8\alpha\phi^{\prime\prime}+1)+B^{\prime}\phi^{\prime}(24\alpha+r^{2}\sigma\phi^{\prime 2}$
$-6\gamma r\phi^{\prime})-((2\gamma+r^{2})(\phi^{\prime})^{2}))-4B^{\prime}(4\alpha\phi^{\prime}$ (45)
$+r)+B^{2}(32\alpha\phi^{\prime\prime}-4\phi^{\prime 2}(\gamma-r^{2}\sigma\phi^{\prime\prime})$
$+\kappa r^{2}\phi^{\prime 4}-16\gamma r\phi^{\prime}\phi^{\prime\prime})+4=0\,,$
$(rr):\quad$ $2A^{\prime}B(r^{2}\sigma B\phi^{\prime 3}+8\alpha(3B-1)\phi^{\prime}-6\gamma rB\phi^{\prime 2}$
$-2r)+A(B^{2}\phi^{\prime 2}(-12\gamma-3\kappa r^{2}\phi^{\prime 2})$
$+8r\sigma\phi^{\prime}+2B((2\gamma+r^{2})\phi^{\prime 2}-2)+4)=0\,,$
(ϕ)::italic-$(\phi):\quad$ $2A(\phi^{\prime}(4Br\left(-B\gamma A^{\prime\prime}+2AB\sigma\phi^{\prime\prime}+A\right)$
$+AB^{\prime}(-6B\gamma+2\gamma+r^{2}))+B\phi^{\prime 2}(B(r^{2}\sigma A^{\prime\prime}$
$-6A\kappa r^{2}\phi^{\prime\prime}+4A\sigma)+6Ar\sigma B^{\prime})$
$+2B(4\alpha(B-1)A^{\prime\prime}+A\phi^{\prime\prime}(-2B\gamma+2\gamma$
$+r^{2}))-AB\kappa r\phi^{\prime 3}(3rB^{\prime}+4B))$
$+AA^{\prime}(B^{\prime}(8\alpha(3B-1)+3Br^{2}\sigma\phi^{\prime 2}$
$-12B\gamma r\phi^{\prime})+2B(\phi^{\prime}(-6B\gamma+2Br^{2}\sigma\phi^{\prime\prime}$
$+2\gamma+r^{2})-B\kappa r^{2}\phi^{\prime 3}-4B\gamma r\phi^{\prime\prime}$
$+4Br\sigma\phi^{\prime 2}))+BA^{\prime 2}(-8\alpha(B-1)$
$-Br^{2}\sigma\phi^{\prime 2}+4B\gamma r\phi^{\prime})=0\,.$
@@
Appendix B Perturbation equations
Here, we present the form of the perturbation equations in the frequency domain. These are the equations that we integrate in our search for the QNMs. In the following equations the functions $\mathcal{F}_{i}$ depend on the background solutions and on the frequency $\omega$. They are rather complicated so we avoid providing their exact form here, but we denote the parameters of the theory upon which they depend.